OpenStudy (anonymous):

Help. Find the length of the diameter of a sphere with a volume of 143.8 miles^3

OpenStudy (anonymous):

is ther like a formula or something :/

OpenStudy (anonymous):

5.86 miles 3.25 miles 6.5 miles 11.72 miles

OpenStudy (anonymous):

The Volume of a sphere = (4/3)*pi*r^3 The volume is given which is 143.8 === Don't Forget that , D (diameter) =2r(radius) Can you continue ?

OpenStudy (anonymous):

6.5 miles then?

OpenStudy (anonymous):

Hmm,Let's See 143.8=4/3 * (3.14)*r^3 r^3=3.25 miles , d = 2r =2(3.25)=6.5 miles So yea !

OpenStudy (anonymous):

Of course.. it all makes sense now.. how could I have not thought of that before.. thank you wise one. Here, have a cookie. (*.*.*)

OpenStudy (anonymous):

Ty for the cookie :)